OpenStudy (anonymous):

Find the equation of a circle given its point in the centre being (1,2)

OpenStudy (anonymous):

hey newton, how do you copy and paste someone else's latex

OpenStudy (anonymous):

(x-1)sqare+(x-2)square=r square

OpenStudy (anonymous):

As far as I know, you can't on here, unless they send it to you before they render it in LaTeX.

OpenStudy (anonymous):

\[(x-1)^2 + (y-2)^2 = r^2 \]

OpenStudy (anonymous):

how annoying

OpenStudy (anonymous):

Then how do we go about solving for r?

OpenStudy (anonymous):

well someone send an email to the coders

OpenStudy (anonymous):

You can't solve for r... a circle with centre 1,2 is not unique.

OpenStudy (anonymous):

Wait couldn't we just use another point close to it, say (1,3) and solve for r?

OpenStudy (anonymous):

r is radius boy it s given ithink

OpenStudy (anonymous):

You could try. It wouldn't work (at all), but you could try.